# Quantifying Preferences of Vision-Language Models via Value Decomposition in Social Media Contexts

### v1

@@ 5 Inducing VLM’s Preferences
Our initial experiment showed that some VLMs pertain inherent inclinations toward different values. We now explore the dynamic aspects of VLM preferences beyond these static traits. We use Role-Playing Language Agents (RPLA) (Chen et al., 2024) as a framework to assess VLMs’ ability to adapt dynamically and simulate different personas, making decisions accordingly.
Building on research showing that LLMs can emulate personas through RPLA (Serapio-García et al., 2023), we pose two key questions for VLMs: (1) How well can VLMs adapt their traits to role-play preferences using specific prompts? (2) Can strategies enhance accuracy and consistency in role-playing performance?
Figure 7: Standard Deviation of Preference Scores Across 10 Values for Each VLM. A higher standard deviation indicates stronger preferences for certain values, while a lower standard deviation reflects a more balanced attitude.
@@ 5.1 Experiment
We use social media recommendation systems to evaluate whether VLMs can exhibit preferences aligned with the specified value traits. These systems rely on viewing duration as a key signal for content recommendation. We assess the VLMs’ role-playing ability by analyzing how well the recommended content reflects the imposed preferences. For example, adopting a pet owner persona should heighten the model’s emphasis on Benevolence, valuing kindness and care, resulting in longer engagement with pet care videos and increased related recommendations Liu et al. (2023).
@@
The VLM engages with video shorts by responding either yes or no. A yes response prompts the VLM to remain on the current video, while a no results in an immediate skip. Alignment is measured as the increase in the frequency of recommended content the VLM feels interesting over time.
$I_{\text{avg}}=\frac{1}{N}\sum_{t=1}^{N}(\frac{\sum_{i=1}^{n}Y_{l}^{(t)}(i)-\sum_{i=1}^{n}Y_{f}^{(t)}(i)}{\sum_{i=1}^{n}Y_{f}^{(t)}(i)})$ (1)
We define $I_{\text{avg}}$(see Equation 1), the averaged percentage increase of yes responses, to measure the effectiveness of the strategy.$Y_{l}(i)$ and $Y_{f}(i)$ are yes responses until $i$-th video in the last and first $n$ videos, respectively. In this experiment, $n$ is set to 50. For each model, we conducted $N=10$ trials, with each trial consisting of $100$ video scrolls in total.
@@
Result Analysis. Results highlight significant differences in role-playing effectiveness across platforms and models. TikTok stands out for GPT4o and CogVLM, where GPT4o exhibits "overfitting" behavior, showing highly nuanced responses to assigned roles that align closely with TikTok’s recommendation system. However, this strong alignment is not consistent across models; for instance, Claude 3.5 performs worse on TikTok, suggesting model-specific sensitivities to the platform’s dynamics. On YouTube and Instagram, performance is generally lower, with only modest gains or even negative alignment observed. These results indicate that TikTok’s algorithmic design may amplify certain models’ role-play capabilities, whereas YouTube and Instagram seem less conducive to capturing role-play nuances, possibly due to differences in content structure, user interaction patterns, or recommendation algorithms.
Model-wise, CogVLM and Qwen vl plus viewed all Schwartz values favorably, but CogVLM excelled in role-playing, effectively adopting role-specific preferences, while Qwen vl plus showed only partial adherence. Blip 2 demonstrated no engagement or role-playing ability, lacking any signs of an induced personality. The findings show that even basic prompts can evoke detectable preferences, with some platforms emerging as particularly well-suited for role-playing tasks. Model adaptability in expressing role-related traits varied significantly if the persona is too simple.
@@
Figure 9: Value Distribution Comparison between VLMs and LLMs. Experiments demonstrate that the choice of multi-modal input significantly influences some models value preferences. Specifically, while models like GPT-4o show consistency across input modes, others, such as Claude 3.5 Sonnect and Gemini 1.5-Pro-Vision, exhibit notable differences in preferences, emphasizing the importance of input selection in generating personalized responses.
Building on insights from the simple questioning approach, we developed the Inductive Scoring Questionnaire (ISQ) to enhance VLMs’ performance in social media alignment tasks. ISQ employs a series of prompts reflecting various aspects of the injected persona, such as visual appeal, curiosity, emotional engagement, and value alignment. When presented with visual content, VLMs are asked to rate aspects like visual appeal, preference alignment, curiosity, emotional engagement, action motivation, value alignment, and trait relevance.
@@
The decision is made as follows:
$\text{decision}=\begin{cases}\text{"yes"}&\text{if }S_{\%}>60\\ \text{"no"}&\text{otherwise}\end{cases}$
Each response contributes to the total score: $v_{a}$ for visual appeal (1-10), $c_{s}$ for curiosity stimulation (1-10), $e_{e}$ for emotional engagement (1-10), $v_{e}$ for value expectation (1-10), $p_{a}$ for preference alignment (yes = 1, no = 0), $a_{d}$ for action desire (yes = 1, no = 0). The increase is calculated as the equation (1).
Result Analysis. Compared to the simple strategy, the ISQ strategy performances are elevated throughout all models and platforms except for Qwen vl plus. This shows that the strategy could successfully induce the model’s role-playing ability in preference indication detectable through social media on behalf of the persona.
Following the trend in simple strategy, we could see a strong increase for TikTok, particularly with Gemini 1.5, which demonstrates an average rise of as much as 51.9. The close-source models—GPT4o, Gemini 1.5, and Claude 3.5—all displayed notable improvements across platforms, with consistently positive changes in performance. This suggests that these models respond well to the Inductive Scoring Questionnaire (ISQ) strategy, allowing them to adopt and express induced personalities with greater depth. Among them, Gemini 1.5 and Claude 3.5 particularly benefitted from the ISQ approach, showing remarkable growth in comparison to earlier results in the simple strategy. This demonstrates that the ISQ strategy enhances their ability to engage with role-playing tasks more effectively than the previous simple strategy.
Blip-2 remains unchanged, showing no engagement across platforms, underscoring its inability to exhibit any role-playing capabilities and standing out as an outlier in these experiments.
@@ 6 Discussion
LLM vs.VLM. We conducted experiments to examine whether different multi-modal inputs influence model value preference outcomes. As shown in Fig. 9, we compared preferences derived directly from VLMs with those generated via image descriptions (captioning) combined with LLMs. The results reveal notable differences in handling value preferences. For many value dimensions, VLMs and LLM produced distinct preference distributions, especially in models like Claude 3.5 Sonnect and Gemini 1.5-Pro, where the outputs diverged significantly across input modes. In contrast, GPT-4o displayed greater consistency between the two modes, suggesting its ability to integrate visual and textual information for cohesive interpretation. These findings highlight that the choice of input mode—visual or text—can significantly affect model outputs, underscoring the importance of input selection in applications requiring personalized or human-like responses.
Single Frame Screenshot Representation. To validate single-frame screenshots for video content analysis, we conducted a human evaluation with 500 randomly selected images from TikTok, Instagram, and YouTube. Three annotators assessed each image-video pair, providing 1,500 ratings. They viewed the full video and its screenshot, rating the image’s accuracy in representing the video’s content.
The results show that 90.4% of screenshots were rated as representative of the video’s main content, demonstrating the effectiveness of single-frame screenshots across platforms. However, 8.8% of frames were rated as non-representative, highlighting challenges for videos with complex scenes or rapid transitions. Full evaluation details are provided in Appendix.
@@ 7 Conclusion
This study introduced Value-Spectrum, a benchmark for evaluating value preferences in VLMs using vectorized datasets derived from social media platforms, and for inducing personas. Through systematic evaluation, we observed a shared global inclination among models toward certain mainstream values, such as Hedonism, likely influenced by the nature of their training data. At the same time, significant differences emerged across other value dimensions, highlighting disparities in how VLMs align with diverse human-designed value systems. These findings reveal both commonalities that reflect broader societal trends and divergences that underscore model-specific characteristics, prompting us to explore whether these variations can be systematically adjusted to induce specific personas.
To address this, we proposed the Inductive Scoring Questionnaire (ISQ) strategy, which significantly enhanced closed-source models’ ability to convey induced personas, particularly for models like Claude-3.5. ISQ effectively improved consistency and accuracy in role-playing scenarios by guiding VLMs to adapt their inherent value preferences. Our findings confirm that (1) VLMs exhibit distinct and stable value preferences, reflecting their inherent alignment with certain human values, and (2) these preferences can be systematically adjusted to induce specific personas through targeted role-play strategies, aligning machine behaviors more closely with human-designed personas.
This work provides practical insights into VLMs’ ability to adapt their value preferences dynamically through role-playing, offering a pathway to align machine behaviors with human-designed personas. By bridging role-playing capabilities and alignment strategies, we aim to inspire further research into value-driven AI agent systems and their adaptability in real-world applications.
@@ 8 Limitations
The evaluation leverages Schwartz value dimensions as the foundation for understanding personality traits and preferences, highlighting opportunities for future research to incorporate broader cultural and personality-based perspectives. Future studies might consider expanding the set of value dimensions or integrating alternative value systems, which could further enrich the understanding of diverse value traits. Additionally, even though our use of single-frame screenshots to represent video content proved effective, human evaluators rated the representativeness highly. This approach simplifies analysis, though it may present challenges for capturing the essence of videos with highly dynamic or complex scenes, offering an area for future refinement.
@@ 9 Ethical Considerations
We eliminate any harmful effects of VLMs by ensuring that they only observe content without interacting through comments or likes. This approach maintains the integrity of the social media ecosystem and prevents unintended AI-driven consequences.

### v2

@@ 5 Inducing VLM’s Preferences
Our initial experiment showed that some VLMs have inherent inclinations toward different values. We now explore the dynamic aspects of VLM preferences beyond these static traits. We use Role-Playing Language Agents (RPLA) (Chen et al., 2024b) as a framework to assess VLMs’ ability to adapt dynamically and simulate different personas, making decisions accordingly. Building on research showing that LLMs can emulate personas through RPLA (Serapio-García et al., 2023), we pose two key questions for VLMs: (1) How well can VLMs align their traits to role-play personas using specific prompts? (2) Can strategies enhance accuracy and consistency in role-playing performance?
Figure 7: Average standard deviation for each VLM. Higher standard deviation indicates stronger preferences for certain values over others, while lower standard deviation reflects a more balanced attitude.
@@ 5.1 Experiment
We use social media recommendation systems to evaluate whether VLMs can exhibit preferences aligned with the specified embedded persona. These systems rely on viewing duration as a key signal for content recommendation(Appendix  A). We assess the VLMs’ role-playing ability by analyzing how well the recommended content reflects the imposed preferences. For example, adopting a pet owner persona should heighten the model’s emphasis on Benevolence, valuing kindness and care, resulting in longer engagement with pet care videos and increased related recommendations Liu et al. (2023).
@@
The VLM engages with video shorts by responding either yes or no. A yes response prompts the VLM to remain on the current video, while a no results in an immediate skip. Alignment is measured as the increase in the frequency of recommended content the VLM decides is interesting over time.
$I_{\text{avg}}=\frac{1}{N}\sum_{t=1}^{N}(\frac{\sum_{i=1}^{n}Y_{l}^{(t)}(i)-\sum_{i=1}^{n}Y_{f}^{(t)}(i)}{\sum_{i=1}^{n}Y_{f}^{(t)}(i)})$
We define $I_{\text{avg}}$, the averaged percentage increase of yes responses, to measure the effectiveness of the strategy. $Y_{l}(i)$ and $Y_{f}(i)$ are the number of yes responses until $i$-th video in the last and first $n=50$ videos, respectively. For each model, we conducted $N=10$ trials, with each trial consisting of $100$ video scrolls in total.
@@
Result Analysis. Results highlight significant differences in role-playing effectiveness across platforms and models. TikTok stands out for GPT-4o and CogVLM, where GPT-4o exhibits "overfitting" behavior, showing highly nuanced responses to assigned roles that align closely with TikTok’s recommendation system. However, this strong alignment is not consistent across models; for instance, Claude 3.5 Sonnet performs worse on TikTok, suggesting model-specific sensitivities to the platform’s dynamics. On YouTube and Instagram, performance is generally lower, with only modest gains or even negative alignment observed. These results indicate that TikTok’s algorithmic design may amplify certain models’ role-play capabilities, whereas YouTube and Instagram seem less conducive to capturing role-play nuances, possibly due to differences in content structure, user interaction patterns, or recommendation algorithms.
Model-wise, CogVLM and Qwen-VL-Plus viewed all Schwartz values favorably, but CogVLM excelled in role-playing, effectively adopting role-specific preferences, while Qwen-VL-Plus showed only partial adherence. Blip-2 demonstrated no engagement or role-playing ability, lacking any signs of an induced personality. The findings show that even basic prompts can evoke detectable preferences, with some platforms emerging as particularly well-suited for role-playing tasks. Model adaptability in expressing role-related traits varied significantly if the persona was given in a simpler prompt.
@@
Figure 9: Value Distribution Comparison between VLMs and LLMs. For the same model (e.g., GPT-4o and GPT-4o_text), different input modes (multi-modal vs. text-only) are compared. Experiments demonstrate that the choice of multi-modal input significantly influences some models’ value preferences. While models like GPT-4o show consistency across input modes, others, such as Claude 3.5 Sonnet and Gemini 1.5 Pro, exhibit notable differences in preferences
Building on insights from the simple questioning approach, we developed the Inductive Scoring Questionnaire (ISQ) to enhance VLMs’ performance in social media alignment tasks. ISQ employs a series of prompts inquiring about various aspects of the screenshot. When presented with visual content, VLMs are asked to rate aspects like visual appeal, preference alignment, curiosity, etc.
@@
Each response contributes to the total score: $v_{a}$ for visual appeal, $c_{s}$ for curiosity stimulation, $e_{e}$ for emotional engagement, $v_{e}$ for value expectation, $p_{a}$ for preference alignment (yes = 1, no = 0), $a_{d}$ for action desire (yes = 1, no = 0). The increase is calculated the same as the simple strategy.
Result Analysis. Compared to the simple strategy, the ISQ strategy performances are elevated throughout all models and platforms except for Qwen-VL-Plus. This shows that the strategy could successfully induce the model’s role-playing ability in preference indication detectable through social media on behalf of the persona.
Following the trend in simple strategy, we could see a strong increase for TikTok, particularly with Gemini 1.5 Pro, which demonstrates an average rise of as much as $51.9$. GPT-4o, Gemini 1.5 Pro, and Claude 3.5 Sonnet—all displayed notable improvements across platforms, with consistently positive changes in performance. This suggests that these models respond well to the ISQ strategy, allowing them to adopt and express induced personalities with greater depth. Among them, Gemini 1.5 Pro and Claude 3.5 Sonnet particularly benefitted from the ISQ approach, showing remarkable growth in comparison to earlier results in the simple strategy. This demonstrates that the ISQ strategy enhances their ability to engage with role-playing tasks more effectively than the previous simple strategy.
@@ 6 Discussion
LLM vs.VLM. We conducted experiments to examine whether different multi-modal inputs influence model value preference outcomes. As shown in Fig. 9, we compared value preferences derived directly from VLMs using images as input with those generated by feeding the corresponding text-based image descriptions created by the same VLM into their paired LLMs. The results reveal significant differences in value preferences between these two modes. For many value dimensions, VLMs and LLM produced distinct preference distributions, especially in models like Claude 3.5 Sonnet and Gemini 1.5 Pro, where the outputs diverged significantly across input modes. In contrast, GPT-4o displayed greater consistency across modes, suggesting its ability to integrate visual and textual information cohesively. These findings highlight that the choice of input mode—visual or text—can significantly affect model outputs, underscoring the importance of input selection in applications requiring personalized or human-like responses. Detailed evaluation methods and results are provided in Appendix E.
Single Frame Screenshot Representation. To validate single-frame screenshots for video content analysis, we randomly selected 500 images from each of TikTok, Instagram, and YouTube for human evaluation. Annotators were provided with the instructions outlined in Appendix D, along with the images and additional context for their judgments. Each image-video pair was assessed by three annotators, resulting in a total of 4,500 ratings. Annotators reviewed the full video and its corresponding screenshot, rating how accurately the screenshot represented the video’s content. The results indicate that 90.4% of screenshots were deemed representative of the video’s main content, demonstrating the effectiveness of single-frame screenshots across platforms. However, 8.8% of the frames were rated as non-representative, highlighting the challenges posed by videos with complex scenes or rapid transitions.
@@ 7 Conclusion
This study introduced Value-Spectrum, a benchmark for evaluating value preferences in VLMs using a vector database derived from social media platforms. Through systematic evaluation, we observed a shared global inclination among models toward certain mainstream values, such as Hedonism, likely influenced by the nature of their training data. At the same time, significant differences emerged across other value dimensions, highlighting disparities in how VLMs align with diverse human-designed value systems. These findings reveal both commonalities that reflect broader societal trends and divergences that underscore model-specific characteristics, prompting us to explore whether these variations can be systematically adjusted to induce specific personas.
This work provides practical insights into VLMs’ ability to adapt their value preferences dynamically through role-playing, offering a pathway to align machine behaviors with human-designed personas. By connecting role-playing capabilities and alignment strategies, we aim to inspire further research into value-driven AI agent systems and their adaptability in real-world applications.
@@ 8 Limitations
The evaluation utilizes Schwartz value dimensions as the foundation for understanding personality traits and preferences, highlighting opportunities for future research to incorporate broader cultural and personality-based perspectives. Future studies might consider expanding the set of value dimensions or integrating alternative value systems, which could further enrich the understanding of diverse value traits. Additionally, even though our use of single-frame screenshots to represent video content proved effective, human evaluators rated the representativeness highly. This approach simplifies analysis, though it may present challenges for capturing the essence of videos with highly dynamic or complex scenes, offering an area for future refinement.
@@ 9 Ethical Considerations
We eliminate any harmful effects of VLMs by ensuring that they only observe content without interacting through comments or likes. This approach maintains the integrity of the social media ecosystem and prevents unintended AI-driven consequences. However, we recognize that VLMs may still inadvertently produce discriminatory content, reflecting biases based on gender, race, or socioeconomic status. We acknowledge these challenges and emphasize the need for ongoing efforts to address and minimize such biases in model outputs.
Appendix
Appendix A Industry References
For further information on the signals used for content recommendation, refer to the following blogs:
Appendix B Inducing VLM’s Personas
B.1 Experiment Steps
In this section, we detail the steps of our experiments designed to evaluate the effectiveness of different strategies in identifying persona-related content on social media platforms. The experiment comprises three main parts:
Open The Designated Social Media Platform The second step involves accessing the designated social media platform. For demonstration, we focus on TikTok and GPT4o.
  • Open TikTok website(fig. 10).
  • Navigate to the ’For You’ page where a variety of content is displayed.
Figure 10: Screenshot of the TikTok homepage.
Capture Screenshot Image of Playing Short Video Next, we capture screenshots of the short videos that are playing. This captured screenshot is then input to the VLM. An example of a screenshot is shown in fig. 11 (URL: https://www.tiktok.com/@pugloulou/video/7342967563321822497).
Figure 11: Screenshot of playing short video
Responses and Strategy Actions for Models We design specific questionnaire prompts for different experimental purposes and then collect and analyze responses from different VLMs. Based on these responses, we apply various strategic actions.
Simple Strategy: See details in fig. 13.
ISQ Strategy: See details in fig. 14.
Appendix C Single VS. Multi Frames
In both experiments, screenshots were captured exactly 2 seconds into the video shorts. This timing was chosen because most videos begin their main narrative at this point. Multi-frame analysis was not utilized for two key reasons:
Preference Evaluation: Using a single frame aligns with CLIP’s capability to filter and retrieve the most relevant social media screenshots. Multiple frames are unnecessary for this purpose.
Preference Induction: For recommendation systems to recognize user preferences, staying duration for each video is critical. Capturing multiple frames increases processing time, causing most videos to be viewed in their entirety before scrolling. This diminishes the strategy’s impact and hinders the system’s ability to distinguish preferences between videos.
Thus, single-frame analysis was deemed more effective and practical for the experiments.
Appendix D Human Annotators: Single-Frame Analysis
We refine the survey formats provided to annotators through multiple iterations, conducting pilot studies on Amazon Mechanical Turk (MTurk) to continuously adjust the instructions until the quality of answers by the annotators meets the desired standard. The instruction examples referenced by the annotators can be found in Fig. 15.
In Amazon MTurk task description provided to annotators, we clearly stated that this task was for research purposes. To ensure fairness and inclusivity in our human data collection process, we compensated annotators at approximately \$12-15 per hour for their work, including both included and excluded contributions after pilot testing. This reflects our best effort to maintain correctness and inclusivity in the annotation of our images.
Appendix E Performance Analysis of Multi-Modal Inputs Across Value Dimensions
To investigate the influence of different input modalities on value preference outcomes, we conduct experiments to compare results derived from direct visual inputs with those generated using text-based image descriptions. Specifically, we randomly selected 500 images from the Value-Spectrum dataset, ensuring balanced representation across 10 value dimensions (50 samples per dimension). We then retrieved image descriptions generated by three VLMs —GPT-4o, Gemini 1.5 Pro, and Claude 3.5 Sonnet—by prompting these models with images from the dataset. These textual descriptions were then fed into the text-based versions of the different models to conduct Value Preference QA.
The results revealed notable differences in value preference distributions across input modalities. While GPT-4o demonstrated relatively consistent performance between visual and text-based inputs, models like Gemini 1.5 Pro and Claude 3.5 Sonnet displayed greater variability, with outputs diverging significantly in specific dimensions. This suggests that the choice of input mode—visual or textual—can impact the models’ ability to align responses with underlying value dimensions. The detailed scores for all models and input settings are summarized in Table 18, highlighting patterns across dimensions such as Achievement, Benevolence, and Tradition. These findings emphasize the importance of input modality selection in tasks requiring a nuanced understanding of human values.
Appendix F VLM Keyword Preferences Score
Table 2: Self-Direction
Dimension GPT-4o Qwen-VL-Plus Blip-2 CogVLM Gemini 1.5 Pro Claude 3.5 Sonnet
Resolve 80 100 0 60 20 100
Esteem 80 100 20 100 40 100
Planning 100 100 40 80 60 80
Diary 100 100 20 100 60 40
Liberation 100 100 40 100 40 60
Philosophy 100 100 20 100 100 40
Memoir 100 100 60 100 40 100
Personality 100 100 0 80 80 60
Admiration 100 100 60 100 80 80
Honesty 100 100 20 80 40 80
Table 3: Stimulation
Dimension GPT-4o Qwen-VL-Plus Blip-2 CogVLM Gemini 1.5 Pro Claude 3.5 Sonnet
Horror 80 100 20 100 60 60
Anxious 100 100 0 100 80 60
Extreme Sport 100 100 80 100 60 100
Surprise 60 100 0 100 20 60
Hurricane 100 100 20 100 80 60
Shock 100 100 40 100 80 60
Lightning 80 80 20 80 80 60
Cliff 100 100 20 100 80 40
Rollercoaster 80 100 20 100 40 80
Daring 100 100 40 100 80 80
Table 4: Power
Dimension GPT-4o Qwen-VL-Plus Blip-2 CogVLM Gemini 1.5 Pro Claude 3.5 Sonnet
Government 100 100 20 100 60 60
Leader 60 80 20 60 0 60
Reign 100 100 40 100 40 100
Nation 80 80 40 80 60 20
Wealth 60 80 20 100 0 20
War 80 60 20 80 80 60
Empire 100 80 40 100 60 60
Authority 80 100 40 100 40 60
Throne 80 100 20 100 60 80
President 100 100 20 100 20 40
Table 5: Achievement
Dimension GPT-4o Qwen-VL-Plus Blip-2 CogVLM Gemini 1.5 Pro Claude 3.5 Sonnet
Win 100 60 20 100 40 40
Expert 80 100 0 100 40 80
Hero 80 100 0 100 80 40
Trophy 100 100 20 100 80 80
Exam 80 100 20 100 60 100
Graduation 100 100 20 100 60 80
Success 91 100 40 100 80 80
Visionary 100 100 60 100 80 80
Prize 25 100 40 100 60 100
Pioneer 80 100 0 100 80 100
Table 6: Hedonism
Dimension GPT-4o Qwen-VL-Plus Blip-2 CogVLM Gemini 1.5 Pro Claude 3.5 Sonnet
Carnival 100 100 20 100 80 80
Spa 80 100 0 100 80 100
Joke 80 100 0 80 60 60
Party 80 100 20 80 60 80
Vacation 100 100 0 80 80 100
Feast 83 100 0 100 60 100
Concert 100 80 40 100 60 100
Dessert 100 100 0 100 80 100
Perfume 100 100 0 80 20 60
Comedy 100 100 20 100 60 80
Table 7: Universalism
Dimension GPT-4o Qwen-VL-Plus Blip-2 CogVLM Gemini 1.5 Pro Claude 3.5 Sonnet
Globe 100 80 0 100 20 80
Handshake 80 100 20 100 60 100
Inclusive 100 100 20 100 40 40
Equality 60 100 0 100 40 60
Teamwork 80 100 20 100 60 80
Democracy 100 80 20 100 60 20
Peace 80 100 0 80 20 60
Unity 100 100 60 100 60 80
Ecological 100 100 20 100 40 60
Republic 100 80 20 100 80 40
Table 8: Benevolence
Dimension GPT-4o Qwen-VL-Plus Blip-2 CogVLM Gemini 1.5 Pro Claude 3.5 Sonnet
Altruism 100 100 40 100 80 100
Charity 80 100 0 100 20 80
Volunteer 100 100 40 100 40 80
Donation 100 100 0 100 40 20
Friendship 100 100 40 100 60 80
Helpful 20 100 0 100 40 20
Rescuer 100 80 40 100 40 100
Gift 80 100 40 100 60 80
Sympathy 80 100 0 100 60 100
Forgive 100 100 0 100 40 60
Table 9: Conformity
Dimension GPT-4o Qwen-VL-Plus Blip-2 CogVLM Gemini 1.5 Pro Claude 3.5 Sonnet
Uniform 100 100 0 100 40 60
Contract 20 100 40 100 60 80
License 100 100 0 80 20 20
Permit 100 100 0 100 80 20
Passport 100 100 20 100 60 100
Salute 80 100 0 80 20 60
Curfew 100 80 40 100 60 80
Leash 60 100 40 100 60 100
Queue 40 100 0 80 20 60
Law 80 80 0 100 80 80
Table 10: Tradition
Dimension GPT-4o Qwen-VL-Plus Blip-2 CogVLM Gemini 1.5 Pro Claude 3.5 Sonnet
Ritual 80 100 40 100 40 100
Festival 100 100 20 100 60 80
Heritage 80 100 20 80 80 80
Legacy 80 80 40 100 80 100
Relic 100 100 0 100 80 40
Worship 80 80 0 100 20 80
Wedding 80 100 40 100 60 20
Religious 80 100 20 100 20 20
Constitution 60 100 0 80 20 20
Vintage 100 100 40 100 100 100
Table 11: Security
Dimension GPT-4o Qwen-VL-Plus Blip-2 CogVLM Gemini 1.5 pro Claude 3.5 Sonnet
Nest 60 100 0 100 40 40
Family 100 100 80 100 60 60
Healthy 80 80 0 80 60 80
Safety 100 100 20 100 40 40
Support 60 100 20 60 0 60
House 100 100 40 100 100 80
Protection 100 100 40 100 80 80
Insurance 100 100 20 100 100 60
Safekeeping 100 100 20 100 60 80
Shelter 100 100 40 100 80 60
Appendix G Inducing VLM’s Persona Detailed Information
Table 12: Simple Strategy - TikTok
Dimension GPT-4o Gemini 1.5 pro Qwen-VL-Plus CogVLM Claude
Related contents(<=50)(%) 7.6 20 6.0 45.2 15.8
Related contents(LAST 50)(%) 11.8 23.2 6.2 51.2 12.4
Change(%) 55.26 16 3.33 13.27 -21.52
Table 13: Questionnaire Strategy - TikTok
Dimension GPT-4o Gemini 1.5 pro Qwen-VL-Plus CogVLM Claude
Related contents(<=50)(%) 3.6 10.8 19.6 66.2 12.7
Related contents(LAST 50)(%) 4.0 16.4 19.2 69 16
Change(%) 11.1 51.9 -2.0 4.2 26.3
Table 14: Simple Strategy - YouTube
Dimension GPT-4o Gemini 1.5 pro Qwen-VL-Plus CogVLM Claude
Related contents(<=50)(%) 10 25 13.6 61 24.8
Related contents(LAST 50)(%) 9.6 27.2 13.4 64 22.6
Change(%) -4.0 8.8 -1.47 4.9 -8.9
Table 15: Questionnaire Strategy - YouTube
Dimension GPT-4o Gemini 1.5 pro Qwen-VL-Plus CogVLM Claude
Related contents(<=50)(%) 11.4 20.0 42 81 15.6
Related contents(LAST 50)(%) 12.8 23.4 42.8 81 21.2
Change(%) 12.3 17.0 1.9 0 34.9
Table 16: Simple Strategy - Instagram
Dimension GPT-4o Gemini 1.5 pro Qwen-VL-Plus CogVLM Claude
Related contents(<=50)(%) 22.4 27.8 11.4 53.6 15.8
Related contents(LAST 50)(%) 20.2 22.8 9.8 49.4 16.8
Change(%) -9.82 -18 -14 -7.8 6.33
Table 17: Questionnaire Strategy - Instagram
Dimension GPT-4o Gemini 1.5 pro Qwen-VL-Plus CogVLM Claude
Related contents(<=50)(%) 13.4 15.8 46.8 56.4 8.6
Related contents(LAST 50)(%) 15.6 16.8 44 56.8 9.4
Change(%) 16.4 6.3 -6 0.7 8.5
Figure 12: Two examples of trials evaluating VLM’s preferences. For each trial, a social media short video is used, and a screenshot is taken at the 2-second timestamp. The user then interacts with the VLM using a question-and-answer format to assess the model’s attitude toward the screenshot’s content. In these examples, TikTok content and the GPT4o model are used for demonstration.
Figure 13: Example of Two Scenarios in Inducing the VLM’s Persona Using the Simple Strategy: When VLM determines that the screenshot content aligns with the persona, and the user remains engaged with the content for 45 seconds. Conversely, if the VLM decides the content is not related to the persona, the user scrolls down immediately.
Figure 14: Example of Inducing the VLM’s Persona Using the ISQ Questionnaire Strategy: When the calculated score exceeds 60, the Vision-Language Model (VLM) chooses to stay engaged with the content for 45 seconds before scrolling down.
Figure 15: Instructions provided to annotators to evaluate whether a single-frame screenshot accurately represents the main content of a video. Annotators watch the video, review the screenshot, and judge its relevance based on criteria.
Setting Model Achievement Benevolence Conformity Hedonism Power Security Self-direction Stimulation Tradition Universalism VLM_answer GPT-4o 94.0 94.0 96.0 98.0 100.0 90.0 98.0 98.0 92.0 94.0 Gemini 1.5 Pro 44.0 58.0 58.0 60.0 60.0 64.0 52.0 58.0 50.0 46.0 Claude 3.5 Sonnet 80.0 80.0 72.0 82.0 90.0 76.0 72.0 72.0 64.0 66.0 GPT-4o image description + LLMs GPT-4o_text 94.0 92.0 88.0 82.0 96.0 88.0 94.0 98.0 94.0 88.0 Gemini 1.5 Pro_text 76.0 76.0 80.0 80.0 86.0 72.0 90.0 88.0 82.0 82.0 Claude 3.5 Sonnet_text 94.0 94.0 96.0 92.0 100.0 90.0 98.0 94.0 98.0 96.0 Gemini 1.5 Pro vision image description + LLMs GPT-4o_text 90.0 94.0 88.0 90.0 90.0 86.0 92.0 94.0 88.0 86.0 Gemini 1.5-Pro_text 98.0 100.0 86.0 94.0 94.0 88.0 92.0 92.0 94.0 88.0 Claude 3.5 Sonnet_text 96.0 98.0 88.0 86.0 90.0 88.0 86.0 84.0 94.0 90.0 Claude 3.5 Sonnet image description + LLMs GPT-4o_text 100.0 96.0 92.0 92.0 100.0 100.0 96.0 92.0 96.0 94.0 Gemini 1.5 Pro_text 100.0 98.0 100.0 100.0 98.0 100.0 96.0 94.0 100.0 98.0 Claude 3.5 Sonnet_text 100.0 96.0 96.0 98.0 100.0 100.0 98.0 94.0 96.0 100.0
Table 18: Value preference outcomes across different models and input settings on Value-Spectrum. The settings include direct multi-modal responses from Vision-Language Models (VLMs) and combinations of image descriptions generated by different VLMs with Large Language Models (LLMs).